### v1

# Unraveling Indirect In-Context Learning Using Influence Functions

Abstract
This work introduces a novel paradigm for generalized In-Context Learning (ICL), termed Indirect In-Context Learning. In Indirect ICL, we explore demonstration selection strategies tailored for two distinct real-world scenarios: Mixture of Tasks and Noisy Demonstrations. We systematically evaluate the effectiveness of Influence Functions (IFs) as a selection tool for these settings, highlighting the potential for IFs to better capture the informativeness of examples within the demonstration pool. For the Mixture of Tasks setting, demonstrations are drawn from 28 diverse tasks, including MMLU, BigBench, StrategyQA, and CommonsenseQA. We demonstrate that combining BertScore-Recall (BSR) with an IF surrogate model can significantly improve performance, leading to average absolute accuracy gains of 0.37% and 1.45% for 3-shot and 5-shot setups when compared to traditional ICL metrics. In the Noisy Demonstrations setting, we examine scenarios where demonstrations might be mislabeled. Our experiments show that reweighting traditional ICL selectors (BSR and Cosine Similarity) with IF-based selectors boosts accuracy by an average of 2.90% for Cosine Similarity and 2.94% for BSR on noisy GLUE benchmarks. In sum, we propose a robust framework for demonstration selection that generalizes beyond traditional ICL, offering valuable insights into the role of IFs for Indirect ICL.
1 Introduction
In-Context Learning (ICL) has emerged as a powerful method for utilizing large language models (LLMs) to handle novel tasks at inference Mann et al. (2020); Min et al. (2022). Unlike traditional approaches that require task-specific fine-tuning, ICL allows a single model to adapt to different tasks without additional training, relying solely on the demonstrations provided in the context. This flexibility not only reduces the cost of task adaptation but also offers a transparent and easily customizable way of guiding the model’s behavior Liu et al. (2021a); Wei et al. (2022). By leveraging the context provided in prompts, ICL has been shown to improve both generalization across diverse tasks and reasoning abilities Anil et al. (2022); Drozdov et al. (2022). Despite its advantages, the success of ICL is closely tied to the choice of demonstrations used in the prompt. Even slight variations in these demonstrations can significantly influence the model’s performance, as shown in numerous studies Zhao et al. (2021); Liu et al. (2021a); Lu et al. (2022).
Figure 1: An example showcasing demonstration selection for Indirect ICL using Influence Functions (IFs). Consider web corpora with many tasks (different from the end-task) and noisy data– Indirect ICL can be formalized as: Mixture of Tasks (Section 3.1) and Noisy (Section 3.2) ICL, respectively. In MoT, for a given target task (e.g. Medical Genetics), we first filter from this (indirect) pool of candidate demonstrations using BertScore and Cosine Similarity, then re-rank with IFs to select suitable demonstrations (e.g. High-School Biology).
Traditional ICL makes numerous assumptions that restrict its applicability to real-world problem domains. For instance, traditional ICL Min et al. (2021); Conneau (2019); Halder et al. (2020) assumes that demonstrations to be selected are directly and accurately annotated for the end-task. However, this is not always the case – for low-resource, sparse, or specialized domains, end-task information and labeled demonstrations might not be available. Similarly, when LLMs are deployed as services, the user query or the end task itself could be unknown beforehand, let alone providing direct demonstrations at inference. Thus, in this paper, we explore a more generalized setting for ICL, which we refer to as Indirect ICL. In indirect ICL, we aim to provide indirect (or incidental) supervision Yin et al. (2023); Li et al. (2024) by selecting demonstrations from a pool of examples where the majority are not directly suited to the end task due to severe distribution shifts. This includes selecting demonstrations from a pool that predominantly consists of examples belonging to other tasks, with few demonstrations from the end task possibly included. Additionally, the demonstration set may be mislabeled by humans Yan et al. (2014); Zhu et al. (2022) or LLMs Wu et al. (2023). Since the effectiveness of ICL heavily relies on the quality of demonstrations selected Kossen et al. (2024); Wu et al. (2022); Wang et al. (2024), selecting the most helpful indirect demonstrations becomes imperative in these situations.
Despite these potential issues with the demonstration set, we wish to pave the way for extracting maximal benefit from any type of annotated dataset, irrespective of label purity or task relatedness. Thus, in order to combat the aforementioned issues with sub-optimal datasets for ICL, we leverage Influence Functions (IFs) Hampel (1974); Cook and Weisberg (1980). IFs offer a formal method for assessing how individual training data points affect model predictions. They have proven effective in a range of downstream machine learning tasks, including mislabeled data detection Koh and Liang (2017); Pruthi et al. (2020), optimal subset selection Feldman and Zhang (2020); Guo et al. (2020); Xia et al. (2024), model interpretation Han et al. (2020); Grosse et al. (2023); Chhabra et al. (2024b), data attribution Bae et al. (2024), data valuation Choe et al. (2024) and analyzing model biases Wang et al. (2019); Kong et al. (2021).
Traditional (direct) ICL methods that use metrics such as BertScore-Recall (BSR; Gupta et al. 2023a) and cosine similarity Reimers (2019) inherently rely on the semantic similarity between demonstrations and test samples. In this paper, we posit that IFs can be a reasonable measure of affinity between the end task and any (indirect) demonstrations. We show that it is practical to use IF to identify candidate demonstrations that represent a close inductive bias with the end that, and utilize this information for highly accurate demonstration selection in the challenging indirect ICL setting. As our experiments and results will demonstrate, this is indeed the case, and we find that IFs can aid in improved performance when simple semantic similarity is insufficient for demonstration selection.
In sum, our work advances ICL demonstration selection and makes the following key contributions and findings:
  • This work formalizes a new and general paradigm for ICL, namely Indirect In-Context Learning, where we benchmark demonstration selection for two distinct and real-world settings: (a) Mixture of Tasks and (b) Noisy Demonstrations.
  • We propose the use of Influence Functions (IFs) as a viable approach for demonstration selection in these generalized ICL settings.
  • For Mixture of Tasks, combining an IF Surrogate model with BertScore-Recall (BSR) can lead to a 0.37 and 1.45 average absolute increase in performance for $k=3$ and $k=5$ shots compared to the best performing traditional ICL metric.
  • For Noisy Demonstrations, we observe that undertaking a weighted average selection using traditional ICL selectors (BSR and Cosine Similarity) and IF based selectors increases the absolute accuracy on average by 2.90% for Cosine and 2.94% for BSR.
2 Preliminaries
We hereby introduce preliminaries of ICL and IF.
2.1 Traditional In-Context Learning
Before we define the more generalized problem of Indirect ICL, we first define traditional ICL.
In-Context learning. ICL allows LLMs to solve test inputs from novel tasks by presenting a few examples of the task in the prompt. Formally, given a set of input $x$ and output $y$ pairs $\{(x_{i},y_{i})\}_{i=1}^{k}$, prompt template $T$, and the test input $x_{\text{test}}$, ICL using an LLM involves prompting it to conditionally generate the test output $y_{\text{test}}$ according to the following distribution:
$y_{\text{test}}\sim P_{\text{LM}}(\cdot\mid T(x_{1},y_{1},\ldots,x_{k},y_{k},x_{\text{test}}))$
Demonstration Selection. In this work we study the problem of selecting $k$ in-context examples from a pool of $N\gg k$ labeled candidates. This is often necessary due to context length limits and cost considerations (Rubin et al., 2021; Gupta et al., 2023a). Formally, the goal is to select a subset $S\subset\{(x_{i},y_{i})\}_{i=1}^{N}$ of size $k$ that maximizes the probability of generating the desired $y_{\text{test}}$ when the LLM is conditioned on $x_{\text{test}}$ and $S$. It is noteworthy that prior studies mainly consider a task-dependent ICL scenario and assume that candidate demonstrations all directly match the end task Min et al. (2021); Conneau (2019); Halder et al. (2020).
2.2 Indirect In-Conext Learning
Now, we describe two scenarios of indirect ICL, one where the candidate pool comprises of demonstrations from various tasks and the other where the demonstrations may have noisy labels.
Mixture of Tasks. Unlike traditional ICL, where candidate demonstrations match the end task at inference, we consider the more generalized indirect ICL setting where the demonstration pool is task-agnostic. In practice, this setting would allow for pooling annotated demonstrations from various accessible tasks. Formally, given a set of input $x$ and output $y$ pairs $\{(x_{i},y_{i})\}_{i=1}^{k}$, where the pairs ${(x_{i},y_{i})}$ may originate from different tasks than the test input $x_{\text{test}}$, the model is prompted to maximize performance across test tasks.
Noisy Demonstrations. To further generalize the problem of indirect ICL, we also consider noisy supervision that is likely existing in the pool of demonstrations. Formally, let $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ represent the training dataset, where $x_{i}\in X$ is the input and $y_{i}\in Y$ is the corresponding binary label. We randomly select 20% of the data points from $D$ and flip their labels.
Once the noisy dataset is generated, we use it for ICL. Formally, given the noisy set of input-output pairs $\{(x_{i},y_{i})\}_{i=1}^{k}$ and a test input $x_{\text{test}}$, the goal is to conditionally generate the test output $y_{\text{test}}$ based on the noisy training data.
2.3 Influence Functions
Here we formally define how we will use IFs to perform Generalized Indirect ICL.
Let the input space be $X$ and the label space be $Y$. The training dataset is denoted as $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in X$ and $y_{i}\in Y$ are the input and label of the $i$-th data point. Given a loss function $\ell$ and a parameter space $\Theta$, the empirical risk minimization problem is defined as:
$\theta^{*}=\arg\min_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},f_{\theta}(x_{i})),$
where $f_{\theta}:X\to Y$ is the model parameterized by $\theta\in\Theta$. The gradient of the loss for the $i$-th data point with respect to a vector $\eta$ is denoted as:
$\nabla_{\eta}\ell_{i}=\nabla_{\eta}\ell(y_{i},f_{\theta}(x_{i})).$
The IF evaluates the effect of individual training data points on the estimation of model parameters Hampel (1974); Cook and Weisberg (1980); Martin and Yohai (1986). It measures the rate at which parameter estimates change when a specific data point is up-weighted.
Specifically, for $k\in[n]$ and $\epsilon\in\mathbb{R}$, we consider the following $\epsilon$-weighted empirical risk minimization problem:
$\theta^{(k)}(\epsilon)=\arg\min_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},f_{\theta}(x_{i}))+\epsilon\ell(y_{k},f_{\theta}(x_{k})).$
Here, the loss function $\ell(y,f_{\theta}(x))$ is assumed to be twice-differentiable and strongly convex in $\theta$ for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$, the empirical risk minimizer (model weights) $\theta^{*}$ is well-defined, and the influence of the $k$-th data point $(x_{k},y_{k})\in D$ on the empirical risk minimizer (model weights) $\theta^{*}$ is defined as the derivative of $\theta^{(k)}(\epsilon)$ at $\epsilon=0$:
$I_{\theta^{*}}(x_{k},y_{k})\coloneqq\frac{d\theta^{(k)}}{d\epsilon}\bigg{|}_{\epsilon=0}$ $=-H(\theta^{*})^{-1}\nabla_{\theta}\ell(y_{k},f_{\theta}(x_{k})).$
where $H(\theta)\coloneqq\nabla_{\theta}^{2}\left(\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},f_{\theta}(x_{i}))\right)$ is the Hessian of the empirical loss.
The IF $I_{\theta^{*}}(x_{k},y_{k})$ on the empirical risk minimizer $\theta^{*}$ is generalized to assess its effect on prediction loss Koh and Liang (2017). Given a validation dataset $D^{\mathcal{V}}:=\{(x^{\mathcal{V}}_{i},y^{\mathcal{V}}_{i})\}_{i=1}^{m}$, the influence of $(x_{k},y_{k})$ on the validation loss is defined as:
$I(x_{k},y_{k})\coloneqq\Bigg{(}$ $\frac{1}{m}\sum_{i=1}^{m}\nabla_{\theta}\ell(y^{\mathcal{V}}_{i},f_{\theta}(x^{\mathcal{V}}_{i}))\bigg{|}_{\theta=\theta^{*}}\Bigg{)}^{\top}$
×$\times I_{\theta^{*}}(x_{k},y_{k}).$
This gives us
$I(x_{k},y_{k})=-\sum_{i=1}^{m}\Big{(}$ $\nabla_{\theta}\ell(y^{\mathcal{V}}_{i},f_{\theta}(x^{\mathcal{V}}_{i}))^{\top}$
$H(\theta^{*})^{-1}\nabla_{\theta}\ell(y_{k},f_{\theta}(x_{k}))\Big{)}.$
The IF $I(x_{k},y_{k})$ provides insight into how a single data point impacts the validation loss. Essentially, it indicates whether $(x_{k},y_{k})$ contributes positively or negatively to the prediction loss. The more positive the influence value, the more it contributes to the loss decreasing, hence it is a beneficial data point to train the model.
Remark. As discussed above, IFs assume convexity of the loss function, which does not hold for LLMs and deep neural networks. Even though the IF formulations we employ in this paper Kwon et al. (2023); Koh and Liang (2017) make this underlying assumption, we find through empirical observations that for indirect ICL, they can work well. Circumventing the convexity assumption in IF is an ongoing area of research Grosse et al. (2023); Chhabra et al. (2024a) and our framework is flexible enough to accommodate any future IF variants.
Recent work has also shown that theoretically, ICL is akin to implicit fine-tuning and that the transformer attention mechanism is actually undertaking (meta) gradient descent Dai et al. (2023). This implies that influence computation can serve as a beneficial tool to assess the usefulness of the demonstration set samples similar to the fine-tuning set.
3 Proposed Approach
In this section, we describe our approach to select demonstrations in both sub tasks.
3.1 Selecting within Mixture of Tasks
In this scenario, we develop influence-based methods for demonstration selection. Specifically, for each validation example, we compute influence values to identify the most impactful examples from a pool of training examples containing a mixture of tasks. Two approaches are employed to calculate these influence scores:
  • A surrogate-model based method, where a lightweight surrogate model such as RoBERTa Liu (2019) is fine-tuned on the candidate demonstrations to compute influence.
  • A pretrained-gradient based method where the samples are passed through the LLM itself. We then compute the IF using the extracted gradients.
Formally, for each validation example $(x_{\text{val}},y_{\text{val}})$, we compute the influence of each training example $(x_{i},y_{i})\in D_{\text{train}}$, where $D_{\text{train}}$ is the set of the training examples. The influence score $I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))$ quantifies the effect of $(x_{i},y_{i})$ on the loss function evaluated at $(x_{\text{val}},y_{\text{val}})$. Using these computed influence values, we select the top $k$ most influential demonstrations.
We compare two versions of computing the IF after extracting the gradient, DataInf Kwon et al. (2023) and TracIn Pruthi et al. (2020). DataInf is based on an easy-to-compute closed-form expression, leading to better computational and memory complexities than other IF computation methods. TracIn traces how the loss on the test point changes during the training process simply using an inner product of training set gradients. Since it does not compute the Hessian matrix, it is faster than DataInf, but at the cost of lower estimation performance.
Additionally, we compare the influence-only methods with well-performing approaches BertScore-Recall (BSR; Gupta et al. 2023a) and Cosine Similarity Reimers (2019). [^1] We use the implementation from Gupta et al. (2023a, b). These methods excel at capturing semantic similarity between validation and training examples. We also compare with a well performing sparse information retrieval algorithm, BM25 Jones et al. (2000).
Lastly, we combine the previously described approaches by implementing a two-stage selection process. First, we perform an initial pruning of the demonstration pool using either BSR or Cosine Similarity. Specifically, for a given number of desired demonstrations $k$, we prune the dataset to select $2k$ candidates from the original set of labeled examples $\{(x_{i},y_{i})\}_{i=1}^{N}$.
We then apply the IF-based methods to re-rank these remaining examples based on their influence on the validation loss. The final selection of $k$ in-context demonstrations is performed by selecting the top $k$ examples from the re-ranked subset.
3.2 Selecting Noisy Demonstrations
In this setting, we utilize IFs to identify noisy samples within the dataset. Formally, let $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ represent the training dataset. First, we employ IFs to prune the dataset by detecting and removing noisy examples, following which the top $k$ in-context demonstrations are selected using either BSR or Cosine Similarity Gupta et al. (2023a) [^2] We will refer to this approach as IF Pruning..
Additionally, we construct approaches that combine the influence values with the BSR or Cosine Similarity scores. To do so, both the influence values and similarity scores are min-max normalized, resulting in scores scaled between 0 and 1. We then reweigh the scores using a linear combination of the normalized values. Let $\alpha$ and $\beta$ represent the weights assigned to the influence values and the similarity scores, respectively, where $\alpha+\beta=1\quad\text{and}\quad 0<\alpha,\beta<1$, the final combined score for each training example is:
$\text{Score}(x_{i},y_{i})=\alpha\cdot I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))+\beta\cdot S(x_{i},y_{i}),$
where $I((x_{i},y_{i}),(x_{\text{val}},y_{\text{val}}))$ is the influence value and $S(x_{i},y_{i})$ represents either the BertScore Zhang et al. (2019) or Cosine Similarity for the training example $(x_{i},y_{i})$. The top $k$ examples with the highest combined scores are selected as demonstrations. [^3] We will refer to this approach as IF Averaging..
In this setting, we compute influence values using our surrogate model approach. In addition to using the IFs from DataInf Kwon et al. (2023), we also conduct influence experiments using the LiSSA method which is an iterative method to compute the inverse Hessian vector product Agarwal et al. (2017); Koh and Liang (2017).
4 Experiments
Here, we expand upon our experimental setup to conduct the experiments and analyze the results.
4.1 Experimental Setup
We discuss our, dataset details and model used to conduct the experiments.
Evaluation Data. For Mixture of Tasks, we collect a generalized pool of examples from different tasks such that the input $x$ and output $y$ pairs $\{(x_{i},y_{i})\}_{i=1}^{k}$ do not necessarily correspond to the same task as the test input $x_{\text{test}}$. The evaluation task pool contains three samples each from 28 different tasks from MMLU Hendrycks et al. (2020), BigBench Srivastava et al. (2022), StrategyQA Geva et al. (2021) and CommonsenseQA Talmor et al. (2018). We evaluate the ICL accuracy, using this train set, on 12 different tasks from MMLU and BigBench.
For Noisy Demonstrations, we employ the noisy dataset framework from Kwon et al. (2023). In their work, the four binary classification GLUE datasets Wang (2018) MRPC, QQP, QNLI, and SST2 are utilized. To simulate a scenario where a portion of the data samples are noisy, 20% of the training data samples are randomly selected and their labels are flipped. We use these noisy datasets as the candidate pool in our experiments and evaluate the ICL accuracy.
Base LLM. In Mixture of Tasks, for $k=3$ shots, we conduct ICL experiments on Llama-2-13b-chat Touvron et al. (2023), Mistral-7b-v0.3 Jiang et al. (2023) and Zephyr-7b-beta Tunstall et al. (2023). For $k=5$ shots we conduct experiments on Llama-2-13b-chat. We extend on the framework designed by Gupta et al. (2023a, b). The temperature is set to 0 for inference. For Noisy ICL, we conduct experiments on Llama-2-13b-chat. All of our experiments run on 8$\times$×NVIDIA RTX 6000 Ada GPUs.
4.2 Method and Baseline Configurations
Here we expand on the methods and baselines we use for our experiments in both settings.
Mixture of Tasks. We construct 4 IF-only methods. 2 based on the Surrogate Model based approach, Sur and 2 based on the Pretrained LLM weights based approach, Pre. We test Data-Inf and TracIn based versions of these approaches, namely, Surrogate Model-DataInf SurD, Surrogate Model-TracIn SurT, Pretrained Model-DataInf PreD and Pretrained Model-TracIn PreT. As mentioned before, SurD and SurT use RoBERTa as the surrogate model, whereas PreD and PreT use Llama2-13b-chat as the pretrained LLM.
[…]
For baselines, we again compare our IF based approaches with BSR, Cosine Similarity and BM25.
Figure 2: Average performance of different demonstration selection methods across 3 LLMs for $k=3$ shots.
4.3 Results on Mixture of Tasks
We present the results on Mixture of Tasks in Figure 2, the full results for multiple number of shots and multiple LLMs can be seen in Appendix A.1. Additional results for the TracIn IF method are provided in Appendix A.2. Results for Pretrained Gradients combined with BertScore and Cosine Similarity are presented in Appendix A.3.
Combining Surrogate Model DataInf with BertScore results in the best performance. As we can clearly observe in Figure 2 and Table 8, Sur[D,BSR] method had the highest average performance across the tasks, in both 3 and 5 shots. This shows the benefit of combining IF with BertScore as performance increased by 0.56 in $k=3$ shots and by 1.52 in $k=5$ shots. The results also show that the maximal benefit of IF methods is gained in combination with the semantic similarity methods. This is due to the fact that IF can leverage the model’s inductive bias to re-rank the retrieved demonstrations effectively, but the initial $2k$ pruning via BSR is critical to shorten the candidate pool to demonstrations that are semantically relevant enough.
Surrogate Models outperform Pretrained Gradients. We see that surrogate models outperformed Pretrained Gradients in demonstration selection for the Mixture of Tasks setting in both the $k=3$ and $k=5$ shots. The fine-tuning of the surrogate model leads it to better capture the test task affinity of the demonstration pool.
DataInf is better than TracIn as an IF method. The speed gains of TracIn come at a cost of performance as the DataInf method of IF computation routinely outperformed TracIn. TracIn likely underperforms because it does not utilize critical second order gradient information since the Hessian $H(\theta^{*})$ is assumed to be the identity matrix. This trend has also been observed in past work on IF methods Chhabra et al. (2024a).
4.4 Results on Noisy ICL
In this section we provide analysis on the aforementioned General Noisy ICL setting. We also conduct two ablations on varying IF $\alpha$ weight in IF Averaging and varying the Noise level in the datasets.
IF Averaging works better than other baselines. Table 1 and Figure 3 clearly show that doing a weighted average between the surrogate model IF and both Cosine and BertScore leads to performance boosts. Atleast one and if not both of the highest performing methods in each of the datasets we tested were from the averaging method. We see that LiSSA and DataInf are similarly effective, with DataInf being more computationally efficient.
Pruning hurts not helps performance. We can see that pruning actively hurts performance as Figure 3 shows that all 3 types of BertScore pruning and all 3 types of Cosine pruning had lower average scores than BertScore and Cosine Similarity. This might be due to the fact that we are removing potentially helpful samples from the demonstration pool, even if they might have noisy labels.
Table 1: ICL Accuracy across MRPC, QNLI, SST2, and QQP datasets using different methods for Noisy ICL, with 20% noise added to the datasets. The top 2 performers for each dataset are in bold.
Method MRPC QNLI SST2 QQP Rand 70.4 69.6 86.2 70.9 BSR 71.3 74.6 80.4 71.4 Cos 72.3 68.2 82.6 73.2 BM25 70.6 67.6 88.0 71.0 PRU-0.1 Rand[Cos] 70.1 68.4 87.0 69.4 Sur[D,Cos] 69.4 69.8 84.0 71.8 Sur[L,Cos] 68.9 68.2 86.8 70.8 Rand[BSR] 71.1 65.2 86.4 68.4 Sur[D,BSR] 70.6 68.0 82.0 71.0 Sur[L,Cos] 70.1 67.4 88.8 68.2 AVG-0.5 Sur[D,Cos] 75.5 74.8 89.8 67.8 Sur[L,Cos] 70.6 75.8 86.4 73.0 Sur[D,BSR] 74.3 69.6 90.6 73.8 Sur[L,BSR] 73.3 73.4 93.6 69.2
Figure 3: Average performance of the baselines across the 4 datasets.
Table 2: Noisy ICL Accuracy with different hyper-parameters for our methods.
Method MRPC QNLI SST2 QQP PRU-0.2 Rand[Cos] 68.1 68.6 86.6 70.0 Sur[D,Cos] 67.7 67.2 86.2 70.0 Sur[L,Cos] 71.3 65.8 86.8 67.2 Rand[BSR] 69.1 72.2 85.4 69.8 Sur[D,BSR] 70.3 67.4 81.8 71.6 Sur[L,BSR] 70.1 66.4 85.2 70.8 PRU-0.3 Rand[Cos] 70.1 68.4 87.0 67.4 Sur[D,Cos] 68.8 69.2 84.6 70.6 Sur[L,Cos] 70.6 68.2 86.6 72.6 Rand[BSR] 70.8 65.8 86.0 70.8 Sur[D,BSR] 70.6 67.4 84.8 72.0 Sur[L,BSR] 69.4 68.2 87.4 68.4 AVG-0.4 Sur[D,Cos] 75.7 71.4 91.6 62.6 Sur[L,Cos] 73.3 67.2 90.6 75.2 Sur[D,BSR] 74.3 68.8 89.8 65.2 Sur[L,BSR] 56.6 72.8 78.0 73.0 AVG-0.6 Sur[D,Cos] 73.5 69.2 87.2 66.2 Sur[L,Cos] 72.1 68.6 84.6 70.8 Sur[D,BSR] 73.5 69.2 94.2 71.8 Sur[L,BSR] 72.3 65.8 94.2 75.4
Figure 4: ICL Accuracy for BertScore and Cosine Similarity with varying influence weight ($\alpha$) for MRPC 0.2
Varying IF $\alpha$ weight. We perform an analysis on the effect of varying alpha on the MRPC dataset with 20% noise added. We found that generally alpha ranges between 0.4 to 0.6 give the most optimal results as Figure 4 indicates. An explanation is that assigning higher weights to IF scores leads to missing out on semantically relevant samples and assigning higher weights to semantic selection methods leads to including several noisy samples.
Varying Noise level. We also test whether our method can perform well on varying noise levels in the dataset. To test this, we create 2 datasets of MRPC with 10% and 30% noise added. As seen in Table 3, in both the cases IF Averaging outperformed other baselines. With the DataInf configuration performing better for the 10% noise dataset and the LiSSA configuration performing better for the 30% noise dataset.
Table 3: MRPC 10% and 30% Noise added results using various methods (top 2 performers in bold).
Method MRPC 0.1 MRPC 0.3 Rand 70.1 70.8 BSR 70.3 70.3 Cos 70.6 70.6 BM25 73.5 73.0 PRU-0.1 Rand[Cos] 68.6 69.9 Sur[D,Cos] 68.4 69.9 Sur[L,Cos] 68.9 68.9 Rand[BSR] 69.1 68.6 Sur[D,BSR] 66.4 69.1 Sur[L,BSR] 66.7 71.8 AVG-0.5 Sur[D,Cos] 75.0 70.3 Sur[L,Cos] 69.9 76.0 Sur[D,BSR] 73.8 70.8 Sur[L,Cos] 72.1 75.7
4.5 Computational Complexity
We present the worst case time complexity for inference of ours and related method in Table 4. As can be observed, our methods are comparable, if not more efficient than the other baselines. Note that the Sur methods require an additional fine-tuning step on a smaller surrogate model before the gradients are extracted, which the Pre methods do not. Furthermore, note that TracIn as an influence method is much faster than Hessian-based approaches (e.g. DataInf) as it assumes that the Hessian is the identity matrix. While this leads to more efficient influence computation, it comes at the cost of lower estimation performance, as our results with TracIn also show.
Table 4: Computational complexity for each test sample at inference where $N$ is #demonstration samples, $p$ is #model parameters, $d$ is embedding size, $K$ is the max length among all candidates, $L$ is the length (in tokens) of the test input, $Z$ is #ngrams
Method Time Complexity BSR $O(NLKd)$ Cos $O(Nd)$ BM25 $O(NZ)$ PreD $O(Np)$ SurD $O(Np)$ PreT $O(Np)$ SurT $O(Np)$ BSR Combined Methods $O(NLKd)+O(Np)$ Cos Combined Methods $O(Nd)+O(Np)$
5 Related Work
In-Context Learning (ICL). Following the scaling of model sizes and learning resources (Mann et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023), LLMs have gained emergent abilities for efficient inference-time adaptation via ICL Mann et al. (2020). However, ICL is critically sensitive to demonstration pool examples (An et al., 2023; Liu et al., 2021a; Zhang et al., 2022) and selection strategies (Rubin et al., 2021; Mavromatis et al., 2023). One line of work studies example scoring and retrieval, utilizing model-agnostic heuristic metrics like perplexity (Gonen et al., 2022), mutual information (Sorensen et al., 2022), semantic similarity (Liu et al., 2021a; Gupta et al., 2023b), etc. to select demonstrations. Another line of work optimizes selection based on empirically verified desirable features a priori, e.g. diversity (Su et al., 2022; Ye et al., 2023), coverage (Gupta et al., 2023a), etc. However, prior work assumes that the demonstration distribution is aligned with task distribution, which is not always the case (Chatterjee et al., 2024). Our work serves as a first to investigate ICL demonstration selection in the task and dataset quality shifts in the ICL settings.
Influence Functions. Influence functions (IFs) comprise a set of methods from robust statistics Hampel (1974); Cook and Weisberg (1982) that have been recently proposed for deep learning data valuation and can provide a conceptual link that traces model performance to samples in the training set. For gradient-based models trained using empirical risk minimization, IFs can be used to approximate sample influence without requiring actual leave-one-out retraining. For deep learning models, the seminal work by Koh and Liang (2017) utilized a Taylor-series approximation and LiSSA optimization (Agarwal et al., 2017) to compute sample influences. Follow-up works such as Representer Point (Yeh et al., 2018) and Hydra (Chen et al., 2021) sought to improve IF performance for deep learning models, constrained to vision applications. More recently, efficient influence estimation methods such as DataInf (Kwon et al., 2023), Arnoldi iteration (Schioppa et al., 2022), and Kronecker-factored approximation curvature (Grosse et al., 2023) have been proposed which can be employed for larger generative language models, such as LLMs. Some other simpler IF approaches simply consider the gradients directly as a measure of influence (Pruthi et al., 2020; Charpiat et al., 2019), followed by some ensemble strategies (Bae et al., 2024; Kim et al., 2024). Recent work has also found that self-influence only on the training set can be a useful measure for detecting sample influence (Bejan et al., 2023; Thakkar et al., 2023).
IFs have been utilized with great success in a number of application scenarios (e.g. classification Chhabra et al. (2024a); Koh and Liang (2017), generative models Kwon et al. (2023); Schioppa et al. (2022); Grosse et al. (2023), active learning Chhabra et al. (2024b); Liu et al. (2021b), etc.). Moreover, while some recent works have considered using influence for selecting direct demonstrations (Nguyen and Wong, 2023; Van et al., 2024), neither of them has consider their effect on inductive bias selection in the indirect ICL setting, which is the focus of our work.
6 Conclusion
We formalize a new paradigm for generalized In-Context Learning, which we call Indirect In-Context Learning. We analyze two different real-world Indirect ICL settings and propose effective demonstration selection strategies for these scenarios. We explore using Influence Functions (IFs) to leverage the informativeness of the samples in the demonstration pool and the models’ task inductive bias. We find that combining a surrogate model-based IF approach with BertScore performs better when there are an overwhelming majority of irrelevant tasks in the candidate pool. We also find that reweighting the surrogate model-based IF scores with traditional metric scores can be helpful in the case where noisy demonstrations are present. Future work will aim to augment the Pretrained Gradient approach by using better/larger LLMs or finetuning the LLMs.
Limitations
The main limitation of Influence Functions is that they are costly to compute, especially for large datasets and LLMs with lots of parameters. This is why we opted to fine-tune a smaller model such as RoBERTa and use pretrained LLMs for our methods. Further performance gains can be attained at the cost of computational speed if fine-tuned LLMs are employed instead. As research on influence estimation methods for LLMs is currently ongoing, faster influence functions developed in the future can also be utilized with our methods for highly efficient and accurate ICL performance.
Appendix
Appendix A Indirect ICL Results for LLM based Influence
A.1 Full Results for Mixture of Tasks
Following are the full results for the Mixture of Tasks setting. For $k=3$ shots in Tables 5, 6, and 7. For $k=5$ shots in Table 8.
Table 5: Performance across different datasets and demonstration selection methods ($k=3$ shots). The datasets are sampled from sub-tasks of the MMLU and BigBench datasets for Llama-2-13b-chat.
Dataset Rand BSR Cos BM25 PreD SurD Sur[D,BSR] Sur[D,Cos] medical-genetics 80.60 87.00 84.00 79.00 76.00 81.00 86.00 82.00 prof-psychology 68.30 70.00 73.00 65.50 65.00 68.50 72.00 68.50 formal-logic 60.16 59.52 60.32 58.73 61.11 59.52 57.14 55.56 moral-disputes 81.00 78.00 79.50 80.50 78.50 76.00 80.50 76.50 public-relations 72.18 79.09 80.91 73.64 71.82 75.45 79.09 79.09 comp-security 76.80 76.00 80.00 76.00 76.00 80.00 76.00 76.00 astronomy 80.26 80.26 78.95 80.92 74.34 79.61 80.26 78.95 abstract-algebra 57.00 58.00 62.00 55.00 57.00 47.00 72.00 72.00 nutrition 75.50 77.50 79.00 78.00 77.50 77.00 79.50 81.50 high-school-biology 76.70 76.50 78.00 76.50 73.50 79.00 80.50 76.50 formal-fallacies 47.25 52.00 50.00 49.50 47.00 56.50 47.50 50.00 tracking-3 40.20 44.00 39.00 45.00 40.00 37.00 43.50 39.00 Average 68.00 69.87 70.39 68.19 66.48 68.04 71.16 69.63
Table 6: Performance across different datasets and demonstration selection methods ($k=3$ shots). The datasets are sampled from sub-tasks of the MMLU and BigBench datasets for Mistral-7b-v3
Dataset Rand BSR Cos BM25 PreD SurD Sur[D,BSR] Sur[D,Cos] medical-genetics 88.25 88.00 91.00 89.00 88.00 86.00 87.00 89.00 prof-psychology 84.50 84.00 84.50 82.00 85.00 83.00 85.50 84.00 formal-logic 66.47 66.67 69.05 65.08 63.49 66.67 68.25 69.05 moral-disputes 87.00 85.00 85.50 87.00 88.5 86.00 87.00 87.50 public-relations 83.41 82.73 81.82 84.55 84.55 83.64 84.55 81.82 comp-security 87.25 83.00 89.00 88.00 90.00 88.00 86.00 90.00 astronomy 87.34 88.82 89.47 88.16 84.87 86.18 86.18 86.18 abstract-algebra 57.75 63.00 60.00 60.00 59.00 50.00 64.00 64.00 nutrition 84.75 86.50 87.50 83.00 83.00 83.50 88.50 87.00 high-school-biology 85.63 84.00 86.50 85.00 87.00 84.00 87.50 87.00 formal-fallacies 49.63 53.50 50.00 53.50 48.00 46.50 52.50 53.50 tracking-3 46.38 49.00 49.00 49.50 46.50 39.00 48.50 45.50 Average 75.70 76.19 76.95 76.23 75.66 73.54 77.13 77.05
Table 7: Performance across different datasets and demonstration selection methods ($k=3$ shots). The datasets are sampled from sub-tasks of the MMLU and BigBench datasets for Zephyr-7b-beta
Dataset Rand BSR Cos BM25 PreD SurD Sur[D,BSR] Sur[D,Cos] medical-genetics 79.50 80.00 77.00 76.00 78.00 82.00 76.00 78.00 prof-psychology 74.50 74.00 74.50 74.50 72.00 74.50 73.00 74.00 formal-logic 69.44 65.87 65.87 65.08 66.67 71.43 65.87 61.11 moral-disputes 77.63 78.50 78.00 76.50 75.00 78.00 78.50 77.00 public-relations 75.00 80.91 72.73 73.64 76.36 75.45 76.36 76.36 comp-security 76.00 73.00 77.00 78.00 75.00 77.00 79.00 79.00 astronomy 79.77 81.58 80.26 80.26 74.34 80.92 79.61 82.24 abstract-algebra 52.00 53.00 53.00 51.00 51.00 50.00 62.00 55.00 nutrition 75.63 77.00 79.50 75.50 74.00 74.50 75.50 77.50 high-school-biology 77.63 81.00 80.50 79.50 78.50 80.00 78.50 78.00 formal-fallacies 49.75 57.50 55.50 56.50 52.50 55.00 51.50 46.50 tracking-3 49.25 49.50 49.50 51.50 52.50 45.00 49.50 49.50 Average 69.68 70.99 70.28 69.83 68.83 70.31 70.45 69.51
Table 8: Performance across different datasets and demonstration selection methods ($k=5$ shots)
Dataset Rand BSR Cos BM25 PreD SurD Sur[D,BSR] Sur[D,Cos] medical-genetics 80.00 86.00 81.00 81.00 84.00 80.00 84.00 83.00 prof-psychology 71.00 71.00 73.50 66.00 70.00 68.50 77.00 71.00 formal-logic 62.70 59.52 57.94 56.35 58.73 68.25 62.70 61.90 moral-disputes 79.50 77.50 81.00 81.00 82.00 81.00 81.50 81.50 public-relations 70.00 78.18 81.82 76.36 70.91 77.82 80.91 78.18 comp-security 75.00 78.00 82.00 77.00 76.00 77.00 81.00 77.00 astronomy 78.95 85.53 82.89 80.92 80.92 82.89 84.87 81.58 abstract-algebra 52.00 63.00 62.00 58.00 63.00 55.00 67.00 65.00 nutrition 71.50 81.00 80.00 78.00 76.00 78.00 79.00 81.00 high-school-biology 73.00 79.00 80.00 76.50 77.00 79.00 81.50 75.50 formal-fallacies 46.50 48.50 48.00 50.00 47.50 43.50 53.00 43.50 tracking-3 36.50 49.00 47.00 46 42.50 52.00 42.00 39.00 Average 66.38 71.35 71.42 68.93 69.04 70.24 72.87 69.84
A.2 TracIn Results
Here we provide results for the TracIn method of Influence Computation for $k=3$ shots in Tables 9, 10, and 11. We also provide results for $k=5$ shots in Table 12.
Table 9: Performance across different datasets and different TracIn Influence methods for $k=3$ shots for Llama2-13b-chat
Dataset PreT SurT Sur[T,BSR] Sur[T,Cos] medical-genetics 77.00 79.00 81.00 85.00 prof-psychology 67.00 66.00 69.50 70.00 formal-logic 56.35 61.11 59.52 59.52 moral-disputes 79.50 76.00 82.00 79.50 public-relations 73.64 72.73 76.36 77.27 comp-security 78.00 74.00 76.00 79.00 astronomy 80.26 75.66 82.24 79.61 abstract-algebra 57.00 61.00 67.00 63.00 nutrition 79.50 79.50 78.50 81.00 high-school-biology 76.50 76.00 79.00 75.00 formal-fallacies 46.50 42.50 47.50 43.00 tracking-3 41.00 35.50 40.00 39.50 Average 67.69 66.58 69.89 69.28
Table 10: Performance across different datasets and different TracIn Influence methods for $k=3$ shots for Mistral-7b-v0.3
Dataset PreT SurT Sur[T,BSR] Sur[T,Cos] medical-genetics 88.00 85.00 87.00 88.00 prof-psychology 83.00 80.00 84.00 86.50 formal-logic 65.08 69.05 65.87 68.25 moral-disputes 87.50 84.50 84.50 89.00 public-relations 80.91 82.73 85.45 81.82 comp-security 87.00 83.00 86.00 87.00 astronomy 86.18 86.18 88.16 90.13 abstract-algebra 61.00 51.00 62.00 57.00 nutrition 85.00 83.00 88.50 86.00 high-school-biology 86.50 84.50 88.00 85.00 formal-fallacies 47.00 52.00 54.50 62.50 tracking-3 44.00 42.00 35.00 38.00 Average 75.10 73.58 75.75 76.60
Table 11: Performance across different datasets and different TracIn Influence methods for $k=3$ shots for Zephyr-7b-beta
Dataset PreT SurT Sur[T,BSR] Sur[T,Cos] medical-genetics 76.00 77.00 74.00 80.00 prof-psychology 72.50 71.50 72.50 72.50 formal-logic 67.46 69.84 67.46 64.29 moral-disputes 77.50 75.50 76.00 77.50 public-relations 76.36 74.55 79.09 72.73 comp-security 77.00 78.00 76.00 75.00 astronomy 76.32 78.95 81.58 80.92 abstract-algebra 50.00 48.00 53.00 52.00 nutrition 75.50 77.00 75.50 78.00 high-school-biology 77.50 75.50 80.50 78.00 formal-fallacies 50.00 41.00 46.00 50.00 tracking-3 50.50 48.00 49.50 42.50 Average 68.89 67.90 69.26 68.62
Table 12: Performance across different datasets and different TracIn Influence methods for $k=5$ shots for Llama2-13b-chat.
Dataset PreT SurT Sur[T,BSR] Sur[T,Cos] medical-genetics 82.00 78.00 83.00 81.00 prof-psychology 69.00 67.50 73.50 73.50 formal-logic 61.90 61.11 57.14 57.14 moral-disputes 81.00 81.00 82.50 81.50 public-relations 70.00 70.00 73.64 78.18 comp-security 75.00 76.00 77.00 76.00 astronomy 82.24 76.32 83.55 78.95 abstract-algebra 53.00 56.00 65.00 64.00 nutrition 77.50 75.50 79.00 79.50 high-school-biology 77.00 74.50 82.50 77.00 formal-fallacies 45.50 47.00 39.50 48.00 tracking-3 40.50 41.50 46.50 36.50 Average 67.87 67.03 70.23 69.27
A.3 Pretrained Gradient Results
Here we provide results for pretrained gradients method of computing IF. These can be found in for $k=3$ shots in Tables 13, 14 and 15 and for $k=5$ shots in Table 16.
Table 13: Performance across different datasets and Pre-training based demonstration selection methods ($k=3$ shots) for Llama2-13b-chat.
Dataset Pre[D,BSR] Pre[D,Cos] Pre[T,BSR] Pre[T,Cos] medical-genetics 85.00 80.00 86.00 84.00 prof-psychology 68.50 73.50 70.50 69.50 formal-logic 55.56 57.14 57.94 54.76 moral-disputes 80.50 80.00 78.50 82.00 public-relations 80.91 77.27 74.55 78.18 comp-security 75.00 79.00 80.00 76.00 astronomy 80.26 76.32 78.95 77.63 abstract-algebra 57.00 56.00 58.00 61.00 nutrition 81.00 80.00 78.00 80.50 high-school-biology 75.50 73.50 80.00 76.00 formal-fallacies 52.50 48.00 46.50 45.50 tracking-3 48.50 47.50 37.00 38.00 Average 70.02 69.02 68.83 68.59
Table 14: Performance across different datasets and Pre-training based demonstration selection methods ($k=3$ shots) for Mistral-7b-v0.3.
Dataset Pre[D,BSR] Pre[D,Cos] Pre[T,BSR] Pre[T,Cos] medical-genetics 86.00 88.00 86.00 88.00 prof-psychology 87.50 83.50 85.00 85.50 formal-logic 64.29 65.87 65.08 65.87 moral-disputes 86.00 85.00 85.00 85.50 public-relations 85.45 80.00 86.36 84.55 comp-security 82.00 88.00 85.00 89.00 astronomy 88.16 90.13 87.50 89.47 abstract-algebra 62.00 59.00 62.00 60.00 nutrition 86.50 84.00 86.50 84.50 high-school-biology 86.00 85.50 86.50 87.00 formal-fallacies 54.50 46.00 51.00 50.00 tracking-3 49.50 48.50 50.00 53.00 Average 76.49 75.29 76.32 76.87
Table 15: Performance across different datasets and Pre-training based demonstration selection methods ($k=3$ shots) for Zephyr-7b-beta.
Dataset Pre[D,BSR] Pre[D,Cos] Pre[T,BSR] Pre[T,Cos] medical-genetics 82.00 77.00 81.00 82.00 prof-psychology 73.00 70.50 74.50 73.00 formal-logic 69.84 67.48 65.08 66.67 moral-disputes 73.00 76.00 75.50 76.50 public-relations 78.18 77.27 76.36 70.00 comp-security 78.00 73.00 75.00 76.00 astronomy 76.97 78.29 77.63 79.61 abstract-algebra 51.00 55.00 58.00 58.00 nutrition 79.00 74.50 76.50 78.50 high-school-biology 78.00 77.00 80.00 78.50 formal-fallacies 54.50 55.50 46.50 54.50 tracking-3 47.00 47.50 49.00 48.50 Average 70.04 69.08 69.59 70.15
Table 16: Performance across different datasets and Pre-training based demonstration selection methods ($k=5$ shots) for Llama2-13b-chat.
Dataset Pre[D,BSR] Pre[D,Cos] Pre[T,BSR] Pre[T,Cos] medical-genetics 83.00 83.00 82.00 83.00 prof-psychology 73.50 74.00 74.00 71.50 formal-logic 60.32 56.35 62.70 62.70 moral-disputes 82.00 82.50 81.00 79.50 public-relations 75.45 75.45 78.18 77.27 comp-security 77.00 77.00 76.00 80.00 astronomy 82.24 81.58 81.58 77.63 abstract-algebra 60.00 59.00 62.00 60.00 nutrition 81.00 78.50 80.50 79.50 high-school-biology 80.50 78.50 79.50 76.00 formal-fallacies 48.50 50.50 49.00 49.00 tracking-3 50.00 46.00 47.50 51.50 Average 71.13 70.20 71.16 70.63
Appendix B Noisy ICL Results for different parameters
Here we provide the results for different noise levels in the dataset and the different hyperparameters we tested in Table 17.
Table 17: MRPC results using various methods and configurations for 10% and 30% Noise.
Method MRPC 0.1 MRPC 0.3 PRU-0.2 Rand[Cos] 68.9 69.1 Sur[D,Cos] 71.1 69.9 Sur[L,Cos] 70.6 72.3 Rand[BSR] 70.1 68.1 Sur[D,BSR] 70.1 70.1 Sur[L,BSR] 71.3 66.7 PRU-0.3 Rand[Cos] 69.9 69.6 Sur[D,Cos] 71.8 69.4 Sur[L,Cos] 71.3 71.3 Rand[BSR] 69.4 67.9 Sur[D,BSR] 70.3 71.1 Sur[L,BSR] 72.1 73.0 AVG-0.4 Sur[D,Cos] 71.8 69.9 Sur[L,Cos] 69.4 73.0 Sur[D,BSR] 72.1 75.4 Sur[L,BSR] 67.4 70.3 AVG-0.6 Sur[D,Cos] 70.3 73.5 Sur[L,Cos] 74.3 73.3 Sur[D,BSR] 70.3 71.8 Sur[L,BSR] 71.3 73.8